OpenStudy (anonymous):

What are the solutions of 3x2 + 6x + 6 = 0?

OpenStudy (dumbcow):

factor out a 3 first

OpenStudy (dumbcow):

--> x^2 +2x+2 = 0

OpenStudy (anonymous):

3 (x^2+2 x+2) = 0

OpenStudy (anonymous):

Use the quadratic formula on that. a = 1 b = 2 c = 2 x = -1-i x = -1+i

OpenStudy (dumbcow):

Then use quadratic formula

OpenStudy (anonymous):

No real solutions.. \[b^{2}-4ac = 36-72 < 0\]

OpenStudy (anonymous):

\[x _{1} =( -2 + \sqrt{4-4*1*2} )/ 2*1 =( -2 + \sqrt{-4} )/2\] \[x _{2} =( -2 - \sqrt{4-4*1*2} )/ 2*1 =( -2 - \sqrt{-4} )/2\]